Determine the tangential and centripetal components of the net force exerted on a car (by the ground) when its speed is 30 m/s , and it has accelerated to this speed from rest in 8.0 s on a curve of radius 400 m . The car's mass is 1250 kg
Answer #1

centripetal force = m*v2 / r = (1250 * 302  ) / 400 = 2812.5 N

tangential Force = m*at  =

m*\frac{\Delta V}{\Delta t} = 1250 * \frac{30}{8} = 4687.5 N